# Increasing the clock speed of a thermodynamic computer by adding noise

### v1

@@ Abstract
I Introduction
In classical computing, the energy scales of even the smallest devices, such as transistors and gates, are large compared to that of the thermal energy, $k_{\rm B}T$. As a result, there is a clear separation of scales between signal and noise, enabling deterministic computation. This determinism comes with an energy cost: classical computers operate far above the limits of thermodynamic efficiency, and require large amounts of power and heat dissipation to ensure their reliability Bennett (1982); Landauer (1991); Ceruzzi (2003). As devices shrink to the nanoscale, the energy associated with computation becomes comparable to that of thermal fluctuations, making computation more energy efficient Frank (2017); Wolpert (2019); Conte et al. (2019). However, the comparable scales of signal and noise makes deterministic computation challenging: efficient time-dependent protocols are required to do even the simplest logic operations on the $k_{\rm B}T$-scale Gingrich et al. (2016); Dago et al. (2021); Proesmans et al. (2020a); Zulkowski and DeWeese (2014); Proesmans et al. (2020b); Blaber and Sivak (2023); Barros et al. (2024).
Fields such as thermodynamic computing view thermal fluctuations as a resource rather than a problem Conte et al. (2019); Hylton (2020). Thermodynamic computing makes use of the tendency of physical systems to evolve toward thermal equilibrium to do computation. In this framework, the thermal bath contributes to computation by providing the fluctuations necessary for state changes, and in some modes of operation the signal is the noise, with the equilibrium fluctuations of the degrees of freedom of the thermodynamic computer being the output of the calculation Aifer et al. (2023); Melanson et al. (2023).
@@ I Introduction
In Section II we demonstrate this clock-acceleration procedure in general, using analytic methods. In Section III we provide a simple numerical illustration of the procedure, using a thermodynamic computer program to invert a $4\times 4$ symmetric positive definite matrix. The equilibration time of even this small system can vary by orders of magnitude as the matrix is changed. This broad distribution of timescales indicates the need for a clock-acceleration procedure, because it would be inconvenient for the run time of a thermodynamic computer program to vary by orders of magnitude as its parameters are varied. We conclude in Section IV.
II Increasing the clock speed of a thermodynamic computer
We consider a thermodynamic computer composed of $N$ units with real-valued scalar amplitudes $S_{i}$. Amplitudes could represent physical distances, if units are realized mechanically Dago et al. (2021), or voltage states, if realized by RLC circuits Aifer et al. (2023); Melanson et al. (2023). Units interact via the potential $V({\bm{S}})$, which typically consists of pairwise couplings between units Aifer et al. (2023); Melanson et al. (2023).
@@ II Increasing the clock speed of a thermodynamic computer
Here $\mu$, the mobility, sets the time constant of the computer. For the thermodynamic computers described in Refs. Aifer et al. (2023); Melanson et al. (2023), which are realized by RLC circuits, $1/\mu\propto RC\sim 1$ microsecond. The combination $k_{\rm B}T$ is the thermal energy scale – $k_{\rm B}$ is Boltzmann’s constant and $T$ is temperature – and $\eta$ is a Gaussian white noise with zero mean, $\langle\eta_{i}(t)\rangle=0$, and covariance $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The Kronecker delta $\delta_{ij}$ indicates independence between different noise components, and the Dirac delta $\delta(t-t^{\prime})$ indicates the absence of time correlations. The noise $\eta$ represents the thermal fluctuations inherent in the system. We consider the interaction $V$ and the noise $\eta$ to specify the program of the thermodynamic computer.
Now say that we have a thermodynamic computer program (1) whose dynamics is slow, and so takes a long time to run (and so to equilibrate, if that is our goal). The most direct way to speed up the computer is to increase the mobility parameter $\mu$, which sets the computer’s time constant. However, for RLC circuits we do not have unlimited freedom to do this, and timescales much shorter than microseconds are difficult to arrange.
An alternative is to consider running an accelerated computer program whose solution is identical to that of the original program but which takes less time to run. We can construct such a program as follows. Rescale the interaction term $V({\bm{S}})$ by a factor $\lambda\geq 1$, and introduce to the system an additional source of Gaussian white noise. If the new noise has variance $\sigma^{2}$ then the new program is described by the equation
@@
where $\langle\xi_{i}(t)\rangle=0$ and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. If we then set
@@
where $\tilde{\mu}\equiv\lambda\mu$. Eq. (5) is Eq. (1) with mobility parameter rescaled by a factor $\lambda\geq 1$ (the correlations of $\eta$ and $\xi$ are the same). Thus the accelerated computer program, (2) and (4), runs the original computer program (1) but faster  [^1] A similar effect could be achieved by rescaling $V\to\lambda V$ and increasing temperature $T\to\lambda T$, but we assume this is less practical than adding an external source of noise.. Another way to see this is to rescale time $t\to\lambda t$ in (1); the result is Eq. (5). The accelerated program runs the original program with the clock speed increased by a factor of $\lambda\geq 1$. The thermodynamic behavior of the system is unchanged.
III Numerical illustration of accelerated matrix inversion
To illustrate this clock-acceleration procedure we consider the matrix inversion program of Refs. Aifer et al. (2023); Melanson et al. (2023). Here the computer’s degrees of freedom possess the pairwise interaction
@@ III Numerical illustration of accelerated matrix inversion
where the sum runs over all distinct pairs of interactions. We choose an interaction matrix $J_{ij}$ that is positive definite and symmetric, and so has all eigenvalues non-negative and $N(N+1)/2$ distinct entries. This interaction is shown schematically in the inset of Fig. 1(a). In thermal equilibrium the probability distribution of the computer’s degrees of freedom is $\rho_{0}({\bm{S}})={\rm e}^{-\beta V({\bm{S}})}/\int{\rm d}{\bm{S}}^{\prime}{\rm e}^{-\beta V({\bm{S}}^{\prime})}$, where $\beta^{-1}\equiv k_{\rm B}T$, and so for the choice (6) we have $\langle S_{i}\rangle_{0}=0$ and
@@
Here the brackets $\langle\cdot\rangle_{0}$ denote a thermal average, and $(J^{-1})_{ij}$ denotes the elements of the inverse of the matrix $J_{ij}$. Thus we can invert the matrix $J_{ij}$ by sampling the thermal fluctuations, and specifically the two-point correlations, of the units ${\bm{S}}$ in thermal equilibrium.
Figure 1: Matrix inversion thermodynamic computer program, following Refs. Aifer et al. (2023); Melanson et al. (2023). (a) Probability distribution of the reciprocal of the smallest eigenvalue for $10^{7}$ $4\times 4$ symmetric positive definite matrices $J$. The values associated with the matrices $J_{1}$ and $J_{2}$ are marked by dots. Inset: schematic of the 4-unit thermodynamic computer used to estimate the inverses of $J_{1}$ and $J_{2}$. (b) Error $E_{1}$ [Eq. (9)] in the estimate for $J_{1}^{-1}$ using the computer program (1) run for time $t_{\rm f}$; here and subsequently we state times in units of $\mu^{-1}$. The program is run $n_{\rm s}=10^{4}$ times, from which the averages $\langle S_{i}S_{j}\rangle$ are calculated. (c) The same for the matrix $J_{2}^{-1}$. Note that the horizontal scales in (b) and (c) are different. (d) and (e): Estimates of the 10 distinct elements of $J_{1}^{-1}$ and $J_{2}^{-1}$, indexed by $k$, for the program times $t_{\rm f}$ indicated. Note that the vertical axes in (d) and (e) have different scales.
We consider the case $N=4$. In Fig. 1(a) we show the probability distribution $P$ of the reciprocal $\alpha_{\rm min}^{-1}$ of the smallest eigenvalue of $10^{7}$ symmetric positive definite $4\times 4$ matrices $J$. The matrices were generated by first creating random symmetric matrices, $B$, with 10 distinct entries, each drawn uniformly from the interval $[0,1)$. We then applied Cholesky factorization to obtain a positive-definite matrix $A=B^{\rm T}B$. To ensure that all pairwise interactions in the thermodynamic computer were of order $k_{\rm B}T$ or larger, we normalized the matrix $A$ by dividing its elements by the magnitude of its smallest entry, yielding the matrix $\beta J$. This normalization step was an arbitrary choice.
The reciprocal of the smallest eigenvalue of the matrix $J$ is a rough estimate of the relaxation time of the resulting thermodynamic computer program. We chose two matrices, $J_{1}$ and $J_{2}$, whose reciprocal smallest eigenvalues are 0.011 and 0.239 respectively, and so we expect the program specified by $J_{2}$ to take at least an order of magnitude longer than that specified by $J_{1}$. We shall see that this is the case (and also that the smallest eigenvalue is not a precise estimate of the equilibration time for two-point correlations).
The matrix $\beta J_{1}$ is
@@
where $\Delta t=10^{-4}\mu^{-1}$ is the time step and $X\sim{\mathcal{N}}(0,1)$ is a Gaussian random number with zero mean and unit variance.
We then reset the units of the computer to zero, and repeat the process, obtaining $n_{\rm s}=10^{4}$ sets of measurements in all. On a single device an alternative would be to run one long trajectory and sample periodically (note that the resetting procedure could be run in parallel, if many devices are available). We compute the average $\langle S_{i}S_{j}\rangle$ of all distinct pairwise products over all samples. If $t_{\rm f}$ is long enough for the computer to come to equilibrium, then $\langle S_{i}S_{j}\rangle\approx\langle S_{i}S_{j}\rangle_{0}$ (for sufficiently large $n_{\rm s}$), and these measurements will allow us to recover the inverse of the relevant matrix via Eq. (7).
As a measure of the error $E$ in the thermodynamic computer’s estimate of the elements $\beta^{-1}(J^{-1})_{ij}$, we use the Frobenius norm
@@
In Fig. 1(c) we show the corresponding plot for the matrix $J_{2}$. In this case, equilibrium is attained for run times $t_{\rm f}\gtrsim 200$ meaning that the program times for inverting the matrices $J_{1}$ and $J_{2}$ differ by about three orders of magnitude. The origin of this difference can be understood from the relative scales of the matrix elements involved: $J_{1}$ encodes a thermodynamic driving force of order 10s of $k_{\rm B}T$ (leading to rapid changes of ${\bm{S}}$), and the associated equilibrium fluctuations of the units ${\bm{S}}$ are on the scale of $10^{-2}$ (which take little time to establish). By contrast, $J_{2}$ encodes a thermodynamic driving force of order $k_{\rm B}T$ (leading to slower changes of ${\bm{S}}$), and some of the associated equilibrium fluctuations are 10s of $k_{\rm B}T$ in magnitude (which take more time to establish).
@@
In Fig. 1(d) and (e) we show the thermodynamic computer’s estimate of the 10 distinct matrix elements (indexed by the variable $k$) of $J_{1}^{-1}$ (panel (d)) and $J_{2}^{-1}$ (panel (e)), for the indicated times, upon collecting $n_{\rm s}=10^{4}$ samples. An accurate estimate of $J_{2}^{-1}$ requires times in excess of 200, or of order 0.1 milliseconds for the clock speed of the thermodynamic computer of Refs. Aifer et al. (2023); Melanson et al. (2023). Given that $10^{4}$ samples are used, the total compute time would be of order a second, for a single $4\times 4$ matrix (scaling estimates Aifer et al. (2023); Melanson et al. (2023) suggest that the characteristic time for matrix inversion using an overdamped thermodynamic computer will grow as the square $N^{2}$ of the matrix dimension $N$, but even matrices of equal size can have substantially different equilibration times).
@@
Figure 2: Accelerated matrix inversion program. (a) Error $E_{2}$ [Eq. (9)] in the estimate of the matrix $J_{2}^{-1}$ using the accelerated computer program (2) and (4), run for time $t_{\rm f}=1$ as a function of the clock-acceleration parameter $\lambda$ (black dashed line). This result is overlaid on the data of Fig. 1(c), derived from the original program (1) run for time $t_{\rm f}$ (green). We collect $n_{\rm s}=10^{4}$ samples for each program. As expected from the comparison of (1) and (5), the accelerated program run for $t_{\rm f}$ is equivalent to the original program run for $\lambda t_{\rm f}$. (b) Exact elements of $J_{2}^{-1}$ and those estimated using the accelerated program run for time $t_{\rm f}$ at two values of $\lambda$. The case $\lambda=1$ is equivalent to the original program.
@@
To speed up the computation of $J_{2}^{-1}$ we run the accelerated program specified by Eqns. (2) and (4). The matrix $J_{2}$ is therefore replaced by the matrix $\lambda J_{2}$, and we introduce an additional source of noise with variance (4). In Fig. 2(a) we show the error $E_{2}$ in the computer’s estimate of the elements of $J_{2}^{-1}$, as a function of the clock acceleration parameter $\lambda$. Here the program time is set to $t_{\rm f}=1$, and $n_{\rm s}=10^{4}$ samples are collected. We overlay this result on that of Fig. 1(c), the error associated with the original program (1) run for time $t_{\rm f}$. As expected from comparison of (1) and (5), these results are similar: the accelerated program run for $t_{\rm f}$ is equivalent to the original program run for $\lambda t_{\rm f}$.
@@
In panel (b) we show the inverse matrix elements extracted from the accelerated program (2) and (4) for two values of $\lambda$. As expected from the comparison of (1) and (5), the elements extracted using $\lambda=1000$ for program time $t_{\rm f}=1$ are as accurate as those extracted using the original program run for time $t_{\rm f}=1000$.
IV Conclusions
Nanoscale computer components must operate in the presence of thermal fluctuations. The premise of thermodynamic computing is to use these fluctuations, as a means of state change and as a form of signal. Here we have proposed a way of accelerating the rate of such computation by adding noise to the computer and scaling its inter-component connections appropriately.

### v3

@@ Abstract
Introduction— In classical computing, the energy scales of even the smallest devices, such as transistors and gates, are large compared to that of the thermal energy, $k_{\rm B}T$. As a result, there is a clear separation of scales between signal and noise, enabling deterministic computation. This determinism comes with an energy cost: classical computers operate far above the limits of thermodynamic efficiency, and require large amounts of power and heat dissipation to ensure their reliability Landauer (1961); Bennett (1982); Ceruzzi (2003). As devices shrink to the nanoscale, the energy associated with computation becomes comparable to that of thermal fluctuations, making computation more energy efficient Frank (2017); Wolpert (2019); Conte et al. (2019). However, the comparable scales of signal and noise makes deterministic computation challenging: efficient time-dependent protocols are required to do even the simplest logic operations on the $k_{\rm B}T$-scale Gingrich et al. (2016); Dago et al. (2021); Proesmans et al. (2020a); Zulkowski and DeWeese (2014); Proesmans et al. (2020b); Wimsatt et al. (2021); Blaber and Sivak (2023); Barros et al. (2024).
The field of thermodynamic computing views thermal fluctuations as a resource rather than a problem Conte et al. (2019); Hylton (2020). Thermodynamic computing makes use of the tendency of physical systems to evolve toward thermal equilibrium to do computation. In this framework, the thermal bath contributes to computation by providing the fluctuations necessary for state changes, and in some modes of operation the signal is the noise, with the equilibrium fluctuations of the degrees of freedom of the thermodynamic computer being the output of the calculation Aifer et al. (2024); Melanson et al. (2025).
Figure 1: Classical digital simulation of a thermodynamic computer program for matrix inversion Aifer et al. (2024); Melanson et al. (2025). (a) Probability distribution of the reciprocal of the smallest eigenvalue for $10^{7}$ $4\times 4$ symmetric positive definite matrices $J$. The values associated with the matrices $J_{1}$ and $J_{2}$ are marked by dots. Inset: schematic of the 4-unit thermodynamic computer used to estimate the inverses of $J_{1}$ and $J_{2}$, comprising 4 units and 10 connections. (b) Error $E_{1}$ [Eq. (9)] in the estimate for $J_{1}^{-1}$ using the computer program (1) run for time $t_{\rm f}$; here and subsequently we state times in units of $\mu^{-1}$. The program is run $n_{\rm s}=10^{4}$ times, from which the averages $\langle S_{i}S_{j}\rangle$ are calculated. (c) The same for the matrix $J_{2}^{-1}$. Note that the horizontal scales in (b) and (c) are different. (d) and (e): Estimates of the 10 distinct elements of $J_{1}^{-1}$ and $J_{2}^{-1}$, indexed by $k$, for the program times $t_{\rm f}$ indicated. Note that the vertical axes in (d) and (e) have different scales.
@@
In what follows we demonstrate this clock-acceleration procedure in general, using analytic methods. We provide a simple numerical illustration of the procedure, using a classical digital simulation of a thermodynamic computer program to invert a $4\times 4$ symmetric positive definite matrix. The equilibration time of even this small system can vary by orders of magnitude as the matrix is changed. This broad distribution of timescales indicates the need for a clock-acceleration procedure, because it would be inconvenient for the run time of a thermodynamic computer program to vary by orders of magnitude as its parameters are varied.
Increasing the clock speed of a thermodynamic computer— We consider a thermodynamic computer composed of $N$ units with real-valued scalar amplitudes $S_{i}$. Amplitudes could represent physical distances, if units are realized mechanically Dago et al. (2021), or voltage states, if realized by RLC circuits Aifer et al. (2024); Melanson et al. (2025), or phases, if realized by Josephson junctions Ray and Crutchfield (2023). The performance of the thermodynamic computer is then set by the characteristic time constants of these devices, roughly milliseconds, microseconds, or nanoseconds, respectively. Our simulation model runs on a conventional computer; when realized in hardware, the thermodynamic computer program runs automatically, driven only by thermal fluctuations Aifer et al. (2024); Melanson et al. (2025). Units interact via the potential $V({\bm{S}})$, which in existing thermodynamic computers consists of pairwise couplings between units Aifer et al. (2024); Melanson et al. (2025).
@@
Here $\mu$, the mobility, sets the time constant of the computer. For the thermodynamic computers described in Refs. Aifer et al. (2024); Melanson et al. (2025), which are realized by RLC circuits, $1/\mu\propto RC\sim 1$ microsecond. The combination $k_{\rm B}T$ is the thermal energy scale – $k_{\rm B}$ is Boltzmann’s constant and $T$ is temperature – and $\eta$ is a Gaussian white noise with zero mean, $\langle\eta_{i}(t)\rangle=0$, and covariance $\langle\eta_{i}(t)\eta_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. The Kronecker delta $\delta_{ij}$ indicates the independence of different noise components, and the Dirac delta $\delta(t-t^{\prime})$ indicates the absence of time correlations. The noise $\eta$ represents the thermal fluctuations inherent to the system. We consider the interaction $V({\bm{S}})$ and the noise $\eta$ to specify the program of the thermodynamic computer.
Now say that we have a thermodynamic computer program (1) whose dynamics is slow, and so takes a long time to run (and to equilibrate, if that is our goal). The most direct way to speed up the computer is to increase the mobility parameter $\mu$, which sets the computer’s time constant. However, for RLC circuits we do not have unlimited freedom to do this, and timescales much shorter than microseconds are difficult to arrange.
An alternative is to consider running an accelerated computer program whose solution is identical to that of the original program but which takes less time to run. We can construct such a program as follows. Rescale the interaction term $V({\bm{S}})$ by a factor $\lambda\geq 1$ – which can be done by uniformly rescaling the computer’s coupling constants, no matter how high-order in ${\bm{S}}$ is the potential – and introduce to the system an additional source of Gaussian white noise. Noise can be injected in a controlled way into thermodynamic computers Melanson et al. (2025), single-molecule experiments  Chupeau et al. (2018), and information engines Saha et al. (2023). If the new noise has variance $\sigma^{2}$ then the new program is described by the equation
@@
where $\langle\xi_{i}(t)\rangle=0$ and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. If we set
@@
where $\tilde{\mu}\equiv\lambda\mu$. Eq. (5) is Eq. (1) with mobility parameter rescaled by a factor $\lambda\geq 1$ (the correlations of $\eta$ and $\xi$ are the same). Thus the accelerated computer program, (2) and (4), runs the original computer program (1) but faster  [^1] A similar effect could be achieved by rescaling $V\to\lambda V$ and increasing temperature $T\to\lambda T$, but we assume scaling temperature is less practical than adding an external source of noise..
Another way to understand this result is to rescale time $t\to\lambda t$ in (1): the result is Eq. (5). Put another way, Eq. (5) can be written as Eq. (1) with a modified definition of time. That is, the thermodynamics of the modified program is exactly the same as the original program: the energy barriers may be higher, but the noise is proportionally larger, and so the modified system evolves across this landscape in exactly the way the original system would, and samples the same equilibrium distribution. The difference is that the modified system has a redefined scale of time, and reaches its destination sooner. In the language of computing, the accelerated program runs the original program with the clock speed increased by a factor of $\lambda\geq 1$.
For this reason the clock-rescaling trick will also work with thermodynamic computers that do calculations out of equilibrium Whitelam and Casert (2024): the dynamical ensembles of (1) and (5) are identical, whether in or out of equilibrium. It will also work with nonlinear interactions $J_{1...N}S_{1}...S_{N}$, because under the rescaling $J_{1...N}\to\lambda J_{1...N}$ (and the addition of noise in the correct proportion) the modified Langevin equation looks like the original Langevin equation with a rescaled notion of time.
Numerical illustration of accelerated matrix inversion— To illustrate this clock-acceleration procedure we consider a classical digital simulation of the matrix inversion program of Refs. Aifer et al. (2024); Melanson et al. (2025). Here the computer’s degrees of freedom possess the bilinear pairwise interaction
@@
where the sum runs over all distinct pairs of interactions. We choose an interaction matrix $J_{ij}$ that is symmetric and positive definite, and so has $N(N+1)/2$ distinct entries and all eigenvalues non-negative. This interaction is shown schematically in the inset of Fig. 1(a). In thermal equilibrium the probability distribution of the computer’s degrees of freedom is $\rho_{0}({\bm{S}})={\rm e}^{-\beta V({\bm{S}})}/\int{\rm d}{\bm{S}}^{\prime}{\rm e}^{-\beta V({\bm{S}}^{\prime})}$, where $\beta^{-1}\equiv k_{\rm B}T$, and so for the choice (6) we have $\langle S_{i}\rangle_{0}=0$ and
@@
Here the brackets $\langle\cdot\rangle_{0}$ denote a thermal average, and $(J^{-1})_{ij}$ denotes the elements of the inverse of the matrix $J$. Thus we can invert the matrix $J$ by sampling the thermal fluctuations, specifically the two-point correlations, of the units ${\bm{S}}$ in thermal equilibrium.
We consider the case $N=4$. In Fig. 1(a) we show the probability distribution $P$ of the reciprocal of the smallest eigenvalue $\alpha_{\rm min}$ for $10^{7}$ symmetric positive-definite $4\times 4$ matrices $J$. These matrices were generated by first constructing random symmetric matrices $B$, where each distinct lower-triangular element (including the diagonal) was drawn from a Gaussian distribution with zero mean and standard deviation 0.1. We then formed the matrix $A=BB^{\rm T}$, which is guaranteed to be symmetric and positive definite. To ensure that all pairwise interaction strengths in the thermodynamic computer were of order $k_{\rm B}T$ or larger, we normalized the resulting matrix by dividing all its elements by the magnitude of the smallest (in absolute value) element of $A$, yielding the matrix $\beta J$. This normalization step was arbitrary and used solely for convenience. The matrix $\beta J$ is guaranteed to be positive definite, and so all its eigenvalues are positive (note that individual elements of the matrix can still be negative).
The reciprocal of the smallest eigenvalue of the matrix $J$ provides a rough estimate of the slowest relaxation timescale of the thermodynamic computer governed by Eq. (1), when $J$ appears in (6). The relaxation of Langevin systems with a quadratic potential can be decomposed into modes corresponding to the eigenvectors of $J$. Each mode relaxes exponentially, with a rate proportional to the corresponding eigenvalue Gardiner (2009). The slowest mode therefore relaxes on a timescale inversely proportional to the smallest eigenvalue of $J$, making its reciprocal a useful proxy for the overall equilibration time. We chose two matrices, $J_{1}$ and $J_{2}$, whose reciprocal smallest eigenvalues are 0.011 and 0.239, respectively. Based on this metric, we expect the program defined by $J_{2}$ to require at least an order of magnitude more time to equilibrate than the one defined by $J_{1}$. We shall see that this expectation is borne out in a qualitative sense, although the smallest eigenvalue is not a precise predictor of the equilibration time for two-point correlations.
The matrix $\beta J_{1}$ is  [^2] For ease of display the matrix elements are shown to not more than 6 significant figures.
@@
where $\Delta t=10^{-4}\mu^{-1}$ is the time step and $X\sim{\mathcal{N}}(0,1)$ is a Gaussian random number with zero mean and unit variance. We set $T=1$ in numerical simulations, but retain $\beta$ in equations for clarity.
We then reset the units of the computer to zero, and repeat the process, obtaining $n_{\rm s}=10^{4}$ sets of measurements in all. On a single device an alternative would be to run one long trajectory and sample periodically (note that the resetting procedure could be run in parallel, if many devices are available). Using all $n_{\rm s}=10^{4}$ samples, we compute the average $\langle S_{i}S_{j}\rangle$ of all distinct pairwise products. If $t_{\rm f}$ is long enough for the computer to come to equilibrium, then $\langle S_{i}S_{j}\rangle\approx\langle S_{i}S_{j}\rangle_{0}$ (for sufficiently large $n_{\rm s}$), and these measurements will allow us to recover the inverse of the relevant matrix via Eq. (7).
As a measure of the error $E$ in the thermodynamic computer’s estimate of the elements $\beta^{-1}(J^{-1})_{ij}$ we use the Frobenius norm
@@
In Fig. 1(c) we show the corresponding plot for the matrix $J_{2}$. In this case, equilibrium is attained for run times $t_{\rm f}\gtrsim 200$, meaning that the run times for inverting the matrices $J_{1}$ and $J_{2}$ differ by about three orders of magnitude. The origin of this difference can be understood from the relative scales of the matrix elements involved: $J_{1}$ encodes a thermodynamic driving force of order 10s of $k_{\rm B}T$ (leading to rapid changes of ${\bm{S}}$), and the associated equilibrium fluctuations of the units ${\bm{S}}$ are on the scale of $10^{-2}$ (which take little time to establish). By contrast, $J_{2}$ encodes a thermodynamic driving force of order $k_{\rm B}T$ (leading to slower changes of ${\bm{S}}$), and some of the associated equilibrium fluctuations are of order 10 in magnitude (which take more time to establish).
@@
In Fig. 1(d) and (e) we show the thermodynamic computer’s estimate of the 10 distinct matrix elements (indexed by the variable $k$) of $J_{1}^{-1}$ (panel (d)) and $J_{2}^{-1}$ (panel (e)), for the indicated times, upon collecting $n_{\rm s}=10^{4}$ samples. An accurate estimate of $J_{2}^{-1}$ requires times in excess of 200, or of order 0.1 milliseconds for the clock speed of the thermodynamic computer of Refs. Aifer et al. (2024); Melanson et al. (2025). Given that $10^{4}$ samples are used, the total compute time would be of order a second, for a single $4\times 4$ matrix (scaling estimates Aifer et al. (2024); Melanson et al. (2025) suggest that the characteristic time for matrix inversion using an overdamped thermodynamic computer will grow as the square $N^{2}$ of the matrix dimension $N$, but matrices of equal size can have substantially different equilibration times).
@@
Figure 2: Accelerated matrix inversion program. (a) Error $E_{2}$ [Eq. (9)] in the estimate of the matrix $J_{2}^{-1}$ using the accelerated computer program (2) and (4) run for time $t_{\rm f}=1$, as a function of the clock-acceleration parameter $\lambda$ (black dashed line). This result is overlaid on the data of Fig. 1(c), derived from the original program (1) run for time $t_{\rm f}$ (green). We collect $n_{\rm s}=10^{4}$ samples for each program. As expected from the comparison of (1) and (5), the accelerated program run for time $t_{\rm f}$ is equivalent to the original program run for time $\lambda t_{\rm f}$. (b) Exact elements of $J_{2}^{-1}$ and those estimated using the accelerated program run for time $t_{\rm f}=1$ at two values of $\lambda$. The case $\lambda=1$ is equivalent to the original program.
@@
To speed up the computation of $J_{2}^{-1}$ we can run the accelerated program specified by Eqns. (2) and (4). The matrix $J_{2}$ is therefore replaced by the matrix $\lambda J_{2}$, and we introduce an additional source of noise with variance (4). In Fig. 2(a) we show the error $E_{2}$ in the computer’s estimate of the elements of $J_{2}^{-1}$, as a function of the clock acceleration parameter $\lambda$. Here the program time is set to $t_{\rm f}=1$, and $n_{\rm s}=10^{4}$ samples are collected. We overlay this result on that of Fig. 1(c), the error associated with the original program (1) run for time $t_{\rm f}$. As expected from comparison of (1) and (5), these results are essentially the same: the accelerated program run for $t_{\rm f}$ is equivalent to the original program run for $\lambda t_{\rm f}$.
@@
In panel (b) we show the inverse matrix elements extracted from the accelerated program (2) and (4) for two values of $\lambda$. As expected from the comparison of (1) and (5), the elements extracted using $\lambda=1000$ (lower panel) for program time $t_{\rm f}=1$ are as accurate as those extracted using the original program run for time $t_{\rm f}=1000$ (right panel of Fig. 1(e)).
Conclusions— As Moore’s law runs its course, classical computing is approaching a local minimum of energy efficiency that is about 3 orders of magnitude removed from the thermodynamic limit of efficiency Hooker (2021); Frank (2017); Conte et al. (2019). Quantum computing may supersede classical computing in the long term, but in the short term faces technical challenges to its large-scale deployment Preskill (2018). Thermodynamic computing aims to bridge the gap between the two paradigms by providing new hardware and algorithms for energy-efficient computing on the $k_{\rm B}T$-scale Conte et al. (2019). Its advantage over classical and quantum computing is that it uses thermal fluctuations rather than attempting to overcome or suppress them, but currently we possess relatively few examples of algorithms that can do computation in this fashion. Those that do exist have some promising characteristics. For instance, the thermodynamic matrix inversion algorithm can be run in a time that scales with matrix size less rapidly than do the corresponding classical algorithms Aifer et al. (2024). However, the prefactors of these run times can vary considerably even between matrices of the same size, reflecting the general phenomenon that arbitrary physical systems can display a wide range of equilibration times.